OpenStudy (thepotatoneedshelp):

Based on the side lengths given (a, b, and c), which triangles are right triangles? A.) a=4, b=6, c= 8 B.) a=6, b=8, c= 10 C.) a=5, b=6, c= square root of 61 d.) a=6, b=9, c=12

OpenStudy (fwizbang):

a right triangle obeys the pythagorean theorem: a^2 + b^2 =c^2

OpenStudy (thepotatoneedshelp):

@fwizbang so how would I set it up??

OpenStudy (fwizbang):

They tell you a, b, and c, so put 'em into the pythagorean theorem and see which work....

OpenStudy (thepotatoneedshelp):

@fwizbang i kinda don't have time to figure it out, can you just tell me the answer?? please

OpenStudy (sshayer):

p^2+q^2=r^2

OpenStudy (thepotatoneedshelp):

@mathmate can you help me with the answer??

OpenStudy (mathmate):

@ThePotatoNeedsHelp http://openstudy.com/study#/updates/575dff78e4b0a532abfa9032

OpenStudy (thepotatoneedshelp):

I can't click on it o-o

OpenStudy (mathmate):

Sorry, if it is a time-limited evaluation, we are not allowed to give help. Please work with what the previous two helpers information. If you take out the calculator and work on it, it will actually take less time! lol

OpenStudy (thepotatoneedshelp):

@mathmate it wasn't a time limit, just trying to finish quickly before bed, lol xD

OpenStudy (mathmate):

Ok, no problem! As I said, there is enough information for you to check. Anybody else would have to do the checking in cases like this. I will resume what they gave. use Pythagoras theorem \(a^2+b^2=c^2\) where a and b are the lengths of the \(shorter\) sides (legs), and c is the longest side (hypotenuse). If the relation is satisfied (i.e. \(a^2+b^2=c^2\)), then the triangle is a right triangle. For example, if you are given a triplet (5,12,13). then you check if \(5^2+12^2=13^2\) Now \(5^2+12^2=25+144=169=13^2\) Thus \(a^2+b^2=c^2\) is satisfied, and the triangle with sides 5,12,13 is a right triangle. Proceed to check each triplet given until you find one that satisfies \(a^2+b^2=c^2\).